### v1

@@ 5 Utilization of LLMs
The application scope of LLMs is extensive and can be practically employed in almost any specialized domain [1; 191; 44; 192; 193]. Following pre-training and fine-tuning, LLMs are primarily utilized by designing suitable prompts for various tasks. Leveraging powerful zero-shot capabilities, many tasks can be directly accomplished by guiding LLMs with straightforward prompts. For more complex tasks that cannot be achieved through simple prompts, a few-shot approach involving in-context learning is employed to guide LLMs in task completion. Additionally, incorporating chain-of-thought [194; 195] prompts in the prompt enhances in-context learning by introducing a reasoning process. The pipeline of the in-context learning and chain-of-thought is shown in Figure 6. In some specialized research directions, obtaining intermediate layer representations of LLMs may be necessary. For instance, in neuroscience studies, embedding representations from the model are used to investigate activation regions of brain functions [196; 197; 198; 199].
@@
Generally, there are several approaches to employing LLMs. The first involves accessing the capabilities of robust proprietary models through open API services, such as utilizing the API provided by ChatGPT [17]. The second approach includes deploying open-source LLMs for local use [7]. The third method entails fine-tuning open-source LLMs to meet specific domain standards [41; 200], enabling their application in a particular field, and subsequently deploying them locally. In Table 5, we have compiled information on various open-source LLMs for reference. Researchers can choose from these open-source LLMs to deploy applications that best suit their needs.
@@ 6 Future Directions and Implications
For researchers in the field of AI, working in isolation is becoming increasingly impractical. The future direction of AI development will intertwine with various industries, necessitating close collaboration with professionals from diverse fields. It is crucial to engage in collaborative efforts, bridging research disciplines, and collectively addressing challenges by combining expertise from different domains. Simultaneously, there is a fresh set of requirements for the comprehensive skills of AI researchers. Training and deploying LLMs necessitate proficiency in managing large-scale data and substantial practical experience in distributed parallel training. This criterion underscores the importance for researchers involved in LLM development to possess substantial engineering capabilities, addressing the challenges inherent in the process. Researchers who are interested in the field of LLMs must either possess engineering skills or adeptly collaborate with engineers to navigate the complexities of model development.
As LLMs find widespread applications in societal life, concerns about ethical issues and societal impact are on a continuous rise. This may involve research and improvements in areas such as managing model biases and controlling the risk of misuse. Considering the paramount importance of privacy and data security, the future development of LLMs might involve more federated learning and decentralized approaches to enhance model performance while safeguarding user privacy. Developers should engage in interdisciplinary collaboration with experts from various fields, including decision-making, legal studies, and sociology, to establish standards and ethical frameworks for the development, deployment, and utilization of LLMs, mitigating potential harmful consequences. In terms of public awareness and education, mandatory awareness training should be implemented before large-scale public deployment and applications. This aims to enhance public understanding of the capabilities and limitations of LLMs, fostering responsible and informed use, especially in industries such as education and journalism.

### v2

@@ 5 Utilization of LLMs
The application scope of LLMs is extensive and can be practically employed in almost any specialized domain [1; 193; 46; 194; 195]. Following pre-training and fine-tuning, LLMs are primarily utilized by designing suitable prompts for various tasks. Leveraging powerful zero-shot capabilities, many tasks can be directly accomplished by guiding LLMs with straightforward prompts. For more complex tasks that cannot be achieved through simple prompts, a few-shot approach involving in-context learning is employed to guide LLMs in task completion. Additionally, incorporating chain-of-thought [196; 197] prompts in the prompt enhances in-context learning by introducing a reasoning process. The pipeline of the in-context learning and chain-of-thought is shown in Figure 6. In some specialized research directions, obtaining intermediate layer representations of LLMs may be necessary. For instance, in neuroscience studies, embedding representations from the model are used to investigate activation regions of brain functions [198; 199; 200; 201].
@@
Generally, there are several approaches to employing LLMs. The first involves accessing the capabilities of robust proprietary models through open API services, such as utilizing the API provided by ChatGPT [19]. The second approach includes deploying open-source LLMs for local use [9]. The third method entails fine-tuning open-source LLMs to meet specific domain standards [43; 202], enabling their application in a particular field, and subsequently deploying them locally. In Table 5, we have compiled information on various open-source LLMs for reference. Researchers can choose from these open-source LLMs to deploy applications that best suit their needs.
@@ 6 Future Directions and Implications
For researchers in the field of AI, working in isolation is becoming increasingly impractical. The future direction of AI development will intertwine with various industries, necessitating close collaboration with professionals from diverse fields. It is crucial to engage in collaborative efforts, bridging research disciplines, and collectively addressing challenges by combining expertise from different domains. Simultaneously, there is a fresh set of requirements for the comprehensive skills of AI researchers. Training and deploying LLMs necessitate proficiency in managing large-scale data and substantial practical experience in distributed parallel training. This criterion underscores the importance for researchers involved in LLM development to possess substantial engineering capabilities, addressing the challenges inherent in the process. Researchers who are interested in the field of LLMs must either possess engineering skills or adeptly collaborate with engineers to navigate the complexities of model development [3].
As LLMs find widespread applications in societal life, concerns about ethical issues and societal impact are on a continuous rise. This may involve research and improvements in areas such as managing model biases and controlling the risk of misuse [4]. Considering the paramount importance of privacy and data security, the future development of LLMs might involve more federated learning and decentralized approaches to enhance model performance while safeguarding user privacy. Developers should engage in interdisciplinary collaboration with experts from various fields, including decision-making, legal studies, and sociology, to establish standards and ethical frameworks for the development, deployment, and utilization of LLMs, mitigating potential harmful consequences. In terms of public awareness and education, mandatory awareness training should be implemented before large-scale public deployment and applications. This aims to enhance public understanding of the capabilities and limitations of LLMs, fostering responsible and informed use, especially in industries such as education and journalism.